OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 4.5, 27, 162, 972 ... an=4.5(2)n-1 an=4.5(6)n-1 an=6(2)n-1 an=6(4.5)n-1

OpenStudy (anonymous):

common ratio = 27 / 4.5 = 6 nth term of a G P is a1 r^(n-1) here a1 = 4.5 and r = 6

OpenStudy (anonymous):

so its 4.5 (6)^(n-1)